OpenStudy (anonymous):

If I know the lengths of two sides of a right triangle, how do I find the third?

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

so say i had the valuse for sides a and b as 2 and 3

OpenStudy (anonymous):

2^2 + 3^2 = c^2

OpenStudy (anonymous):

4+9=c^2

OpenStudy (anonymous):

sqrt of 13 = c

OpenStudy (anonymous):

oh alright, Thanks!

OpenStudy (anonymous):

np